# Hybrid Feedback for Three-dimensional Convex Obstacle Avoidance (Extended version)

### v4

@@ V STABILITY ANALYSIS
    the number of jumps is finite.
@@
See Appendix IX-D. ∎
VI IMPLEMENTATION PROCEDURE
@@
We consider a workspace with convex obstacles that satisfies Assumption 8 with some $\delta>0$. The target location is set at the origin within the interior of the obstacle-free workspace $\mathcal{W}_{r_{a}}^{\circ}$. The center of the robot is initialized in the obstacle-free workspace $\mathcal{W}_{r_{a}}$. The variables $\mathbf{h},\mathbf{a},m$ and $s$ are initialized in the sets $\mathcal{W}_{r_{a}},\mathbb{S}^{2},\mathbb{M}$ and $\mathbb{R}_{\geq 0}$, respectively. Since $\gamma\in(0,\delta-r_{s})$, the algorithm to avoid one obstacle at a time, as discussed in Remark 3. The parameters $\gamma_{a}$ and $\gamma_{s}$ are set to satisfy $0<\gamma_{a}<\gamma_{s}<\gamma.$ A sufficiently small value for $\bar{\epsilon}$ is selected as stated in Lemma 1, and the parameter $\epsilon\in(0,\bar{\epsilon}].$ We set a sensing radius $R_{s}>r_{a}+\gamma$ such that the robot can detect the boundary of the obstacles within its line-of-sight inside the region $\mathcal{B}_{R_{s}}(\mathbf{x})$.
@@
When the control input is initialized in the move-to-target mode, according to (11a), it steers the robot straight towards the origin. The robot should constantly measure the distance between its center and the surrounding obstacles to identify whether the state $\xi$ has entered in the jump set $\mathcal{J}_{0}$ of the move-to-target mode. To do this, the robot needs to identify the set $\partial\mathcal{O}$ which contains the locations from the boundary of the surrounding obstacles that are less than $R_{s}$ units away from the center of the robot and have clear line-of-sight to the center of the robot, where $R_{s}>r_{a}+\gamma$ represents sensing radius. In other words, the set $\partial\mathcal{O}$ is defined as
$\partial\mathcal{O}=\{\mathbf{p}\in\partial\mathcal{O}_{\mathcal{W}}|\lVert\mathbf{x}-\mathbf{p}\rVert\leq R_{s},\mathcal{L}_{s}(\mathbf{x},\mathbf{p})\cap\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{\mathcal{W}})=\emptyset\}.$ (33)
@@
Then, one can obtain the distance between the robot’s center and the surrounding obstacles by evaluating $d(\mathbf{x},\partial\mathcal{O})$ according to Section II-B. If $d(\mathbf{x},\partial\mathcal{O})\leq r_{a}+\gamma_{s}$, the algorithm should identify whether the robot can move straight towards the target location without colliding with the nearest obstacle. In other words, the algorithm should evaluate whether the center of the robot belongs to the landing region $\mathcal{R}_{l}$, defined in (15), associated with the nearest obstacle. To that end, the algorithm identifies the set $\partial\mathcal{O}_{c}\subset\partial\mathcal{O}$ which contains the locations from the set $\partial\mathcal{O}$ that belong to the boundary of the closest obstacle. Let $\mathcal{O}_{i}$ for some $i\in\mathbb{I}$ be the closest obstacle to the center of the robot, then the set $\partial\mathcal{O}_{c}$ is defined as
$\partial\mathcal{O}_{c}=\partial\mathcal{O}_{i}\cap\partial\mathcal{O}.$ (34)
Once the set $\partial\mathcal{O}_{c}$ has been identified, one needs to determine whether the center of the robot $\mathbf{x}$ belongs to the landing region $\mathcal{R}_{l}$ by evaluating the intersection between the set $\partial\mathcal{O}_{c}$ and the set $\mathcal{D}_{r_{a}}^{\circ}(\mathcal{L}_{s}(\mathbf{x},\mathbf{0}))$. If $\partial\mathcal{O}_{c}\cap\mathcal{D}_{r_{a}}^{\circ}(\mathcal{L}_{s}(\mathbf{x},\mathbf{0}))\not=\emptyset$, then the center of the robot belongs to the set $\mathcal{R}_{l}$ and, as per (18) and (20), the state $\xi$ has entered in the jump set of the move-to-target mode $\mathcal{J}_{0}$. Otherwise, the robot continues to operate in the move-to-target mode.
@@
When the state $\xi$ enters in the jump set of the move-to-target mode $\mathcal{J}_{0}$, the algorithm updates the state $\xi$ as per (27) and (31), and the control input switches to the obstacle-avoidance mode. According to Lemma 4, when the robot operates in the obstacle-avoidance mode, it stays inside the $\gamma-$neighborhood of the closest obstacle. As the robot operates in the obstacle-avoidance mode, the algorithm continuously evaluates the intersection $\partial\mathcal{O}_{c}\cap\mathcal{D}_{r_{a}}^{\circ}(\mathcal{L}_{s}(\mathbf{x},\mathbf{0}))$ to check whether the center of the robot has entered in the exit region $\mathcal{R}_{e}$. If $\partial\mathcal{O}_{c}\cap\mathcal{D}_{r_{a}}^{\circ}(\mathcal{L}_{s}(\mathbf{x},\mathbf{0}))=\emptyset$, then the center of the robot belongs to the exit region $\mathcal{R}_{e}$. Additionally, if the target location is at least $\epsilon$ units closer to $\mathbf{x}$ than to the current hit point $\mathbf{h}$, then it implies that $\xi\in\mathcal{J}_{1}$. Then the algorithm updates the value of the variables $m$ and $s$ as per (30) and the control input switches to the move-to-target mode.
@@
Finally, if the control input is initialized in the obstacle-avoidance mode, according to (24), the state $\xi(0,0)$ belongs to the jump set of the obstacle-avoidance mode $\mathcal{J}_{1}$. As a result, according to (30) the algorithm updates the value of the variables $\mathbf{h},\mathbf{a},m$ and $s$, and switches the control input to the move-to-target mode.
The above-mentioned implementation procedure is summarised in Algorithm 1.
@@
2:  Initialize $\mathbf{x}(0,0)\in\mathcal{W}_{r_{a}}$, $\mathbf{h}(0,0)\in\mathcal{W}_{r_{a}}$, $\mathbf{a}(0,0)\in\mathbb{S}^{2}$, $m(0,0)\in\mathbb{M}$ and $s(0,0)\in\mathbb{R}_{\geq 0}$. Choose sufficiently small value of $\bar{\epsilon}$ according to Lemma 1, and initialize $\epsilon\in(0,\bar{\epsilon}].$ Identify the parameter $\delta$, as discussed in Assumption 8, and set parameters $\gamma,\gamma_{s}$ and $\gamma_{a}$ such that $0<\gamma_{a}<\gamma_{s}<\gamma<\delta-r_{s}$. Choose $R_{s}>r_{a}+\gamma$, used in (33). Set $\tau_{s}>0$ and $\delta_{s}\in(0,\tau_{s})$.
3:  Measure $\mathbf{x}$ and the set $\partial\mathcal{O}$ as defined in (33).
@@
5:     if $d(\mathbf{x},\partial\mathcal{O})\leq r_{a}+\gamma_{s}$then
6:        Identify the set $\partial\mathcal{O}_{c}\subset\partial\mathcal{O}$ as defined in (34).
7:        if $\partial\mathcal{O}_{c}\cap\mathcal{D}_{r_{a}}^{\circ}(\mathcal{L}_{s}(\mathbf{x},\mathbf{0}))\neq\emptyset$then
@@
16:        if $d(\mathbf{x},\partial\mathcal{O})\leq r_{a}+{\gamma},$ then
17:           Identify the set $\partial\mathcal{O}_{c}\subset\partial\mathcal{O}$ as defined in (34).
18:           if $\partial\mathcal{O}_{c}\cap\mathcal{D}_{r_{a}}^{\circ}(\mathcal{L}_{s}(\mathbf{x},\mathbf{0}))=\emptyset$then
@@
VII SIMULATION RESULTS
We compare our approach with the separating hyperplane method developed in [12] which can be implemented in a priori unknown environments using information from a range-bearing sensor mounted on the robot. However, unlike our approach, it only works for strongly convex obstacles [12, Assumption 2]. The workspace contains two obstacles, an ellipsoidal obstacle (strongly convex) and a polygonal obstacle that does not satisfy [12, Assumption 2]. The robot’s radius is set to $0.2m$, with a minimum safety distance $r_{s}=0.1m$ and $\gamma=0.5m$. The gains $\kappa_{s}=0.4$ and $\kappa_{r}=2$ are used in (11a), while the sensing radius $R_{s}=2m$ is used in (33), and $\epsilon=0.5m$ in (22).
In Fig. 2a, the dotted portion of the trajectory represents the motion of the robot in the move-to-target mode, whereas the solid portion corresponds to the motion in the obstacle-avoidance mode. The proposed hybrid feedback controller guarantees safe, global asymptotic convergence to the target location at the origin starting from eight different initial locations as shown in Fig. 2a. In contrast, the separating hyperplane approach generates trajectories that converge to an undesired equilibrium (local minimum), as seen in Fig. 2b. The complete simulation video is available at222[Online]. Available: https://youtu.be/xihWpec8kis.

### v5

@@ VI Application to sphere worlds
    the number of jumps is finite,
  4. iv)
    the distance from the robot’s location to the target location is monotonically decreasing.
@@
See Appendix -G. ∎
VII Implementation Procedure
@@
We consider a workspace with convex obstacles that satisfies Assumption 9 with some $\delta>0$, as discussed in Section III. The target location is set at the origin within the interior of the obstacle-free workspace $\mathcal{W}_{r_{a}}^{\circ}$. The parameters $\gamma,\gamma_{s}$ and $\gamma_{a}$ are chosen to satisfy $0<\gamma_{a}<\gamma_{s}<\gamma<(\delta-r_{s})$. A sufficiently small value for $\bar{\epsilon}$ is selected, and the parameter $\epsilon$, used in (28), is chosen such that $\epsilon\in(0,\bar{\epsilon}]$. The parameter $\delta_{s}$, used in (31), is set such that $\delta_{s}\in(0,\tau_{s})$, where $\tau_{s}>0$. The robot is equipped with a range scanner with a sensing radius of $R_{s}>r_{a}+\gamma$. The composite state vector $\xi$ is initialized in the set $\mathcal{K}$.
@@
Switching from the move-to-target mode to the obstacle-avoidance mode: When the control input is initialized in the move-to-target mode, according to (12a), it steers the robot straight towards the origin. The robot should constantly measure the distance between its center and the surrounding obstacles to identify whether $\xi$ has entered in the jump set $\mathcal{J}_{0}$ of the move-to-target mode. To do this, the robot needs to identify the set ð$\eth_{\mathcal{W}}^{\mathbf{x}}$, as defined in (8), which contains the locations from the boundary of the surrounding obstacles that are less than $R_{s}$ units away from the center of the robot and have a clear line of sight to the center of the robot, where $R_{s}>r_{a}+\gamma$ represents the sensing radius. Then, one can obtain the distance between the robot’s center and the surrounding obstacles by evaluating $d(\mathbf{x},\eth_{\mathcal{W}}^{\mathbf{x}})$ according to Section II-B.
@@
If $d(\mathbf{x},\eth_{\mathcal{W}}^{\mathbf{x}})\leq r_{a}+\gamma_{s}$, one should identify whether the robot can move straight towards the target location without colliding with the nearest obstacle within the sensing region. In other words, one should check condition (18) to identify whether the center of the robot belongs to the avoidance region $\mathcal{R}_{a}$, defined in (19), associated with the nearest obstacle, let us say $\mathcal{O}_{i},i\in\mathbb{I}$. To that end, one needs to identify the set ð$\eth_{i}^{\mathbf{x}}$, as defined in (17), which contains the locations from the set ð$\eth_{\mathcal{W}}^{\mathbf{x}}$ that belong to the boundary of the closest obstacle $\mathcal{O}_{i}$. Once the set ð$\eth_{i}^{\mathbf{x}}$ has been identified, one needs to determine whether $\mathbf{x}$ belongs to the avoidance region $\mathcal{R}_{a}$, defined in (19), by evaluating the intersection between the set ð$\eth_{i}^{\mathbf{x}}$ and the set $\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})$. If ð$\eth_{i}^{\mathbf{x}}\cap\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})\not=\emptyset$, then the center of the robot belongs to the avoidance region $\mathcal{R}_{a}$ and the state $\xi$ has entered in the jump set $\mathcal{J}_{0}$ of the move-to-target. Otherwise, the robot continues to operate in the move-to-target mode.
@@
Switching from the obstacle-avoidance mode to the move-to-target mode: When the state $\xi$ enters in the jump set $\mathcal{J}_{0}$, the state $\xi$ is updated as per (34) and (38), and the control input switches to the obstacle-avoidance mode. According to Lemma 6, when the robot operates in the obstacle-avoidance mode, it stays inside the $\gamma-$neighborhood of the closest obstacle. As the robot operates in the obstacle-avoidance mode, we continuously evaluate the intersection ð$\eth_{i}^{\mathbf{x}}\cap\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})$ to check whether the center of the robot has entered in the exit region $\mathcal{R}_{e}$. If ð$\eth_{i}^{\mathbf{x}}\cap\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})=\emptyset$, then the center of the robot belongs to the exit region $\mathcal{R}_{e}$. Additionally, if the target location is at least $\epsilon$ units closer to $\mathbf{x}$ than to the current hit point $\mathbf{h}$, it implies that $\xi\in\mathcal{J}_{1}$. Then, the value of the variables $m$ and $s$ are updated as per (37) and the control input switches to the move-to-target mode.
@@
Finally, if the control input is initialized in the obstacle-avoidance mode, according to (31), the state $\xi(0,0)$ belongs to the jump set of the obstacle-avoidance mode $\mathcal{J}_{1}$. As a result, according to (37), the variables $\mathbf{h},\mathbf{a},m$ and $s$, are updated and the control input switches to the move-to-target mode.
The above-mentioned implementation procedure is summarized in Algorithm 1.
VII-A Control-input smoothing mechanism for practical implementations
Since the proposed hybrid closed-loop system (38) satisfies the hybrid basic conditions, as mentioned in Lemma 3, the resultant robot trajectories during the flow are always smooth. However, the control input vector $\mathbf{u}$, defined in (12a), in general, changes direction instantaneously when switching between modes, causing the overall robot trajectories to become non-smooth. This instantaneous change in the direction of $\mathbf{u}$ is caused by a discrete change in the value of the mode indicator variable $m$ when the state $\xi$ enters in the jump set $\mathcal{J}$. Therefore, in order to generate smooth robot trajectories, one solution consists in replacing $m$ in (12a) with the following continuous scalar function $\lambda$ which takes values between 00 and $1$:
$\lambda(m,s,\tau)=\begin{cases}\varphi_{t_{s}}(s-\tau),&m=1,\\ \varphi_{t_{s}}(\tau-s),&m=0,\end{cases}$ (45)
where the scalar variable $\tau$ is kept constant during the flow and is updated only when $\xi$ enters in the jump set $\mathcal{J}$. The update law for $\tau$, which is denoted by $\mathbf{U}(\xi)$ is given by
$\mathbf{U}(\xi)=\begin{cases}s+\tau_{s},&\xi\in\mathcal{J}_{sm}^{0},\\ s+\tau_{s}+t_{s},&\xi\in\mathcal{J}_{sm}^{1},\end{cases}$ (46)
with $t_{s}>0$ and $\tau_{s}>0$ as defined in (34). For $z\in\{0,1\}$, the jump set $\mathcal{J}_{sm}^{z}$ is defined as
$\mathcal{J}_{sm}^{z}:=\{\xi\in\mathcal{J}_{z}\mid s-\tau\geq t_{s}\},$ (47)
where $\mathcal{J}_{z}$ for $z=0$ and $z=1$ is given by (26) and (31), respectively. According to (46) and (47), $\tau$ is updated only when $\xi$ enters in the jump set $\mathcal{J}_{sm}$ and the difference $s-\tau$ is greater than or equal to $t_{s}$ seconds, where $\mathcal{J}_{sm}=\mathcal{J}_{sm}^{0}\cup\mathcal{J}_{sm}^{1}$. This ensures that the consecutive switching instances are separated by at least $t_{s}$ seconds. According to (46) and (47), the variable $\tau$ is updated only when $\xi$ enters the jump set $\mathcal{J}$ and the condition $s-\tau\geq t_{s}$ is satisfied. This ensures that consecutive switching events are separated by at least $t_{s}$ seconds.
Given a scalar variable $p$, the continuous scalar function $\varphi_{t_{s}}(p)$ is defined by
$\varphi_{t_{s}}(p)=\begin{cases}1,&p\geq t_{s},\\ p/t_{s},&0\leq p\leq t_{s},\\ 0,&p\leq 0.\end{cases}$ (48)
Depending on the value of $p$, the function $\varphi_{t_{s}}(p)$ takes values between 00 and $1$.
Notice that according to (34) and (46), when $\xi\in\mathcal{J}_{0}$, the variables $s$ and $\tau$ are updated to $s+\tau_{s}$. Since $\tau$ is kept constant during the flow, one has $\dot{\tau}=0$. Additionally, as per (38), we have $\dot{s}=1$. Therefore, as the robot enters in the obstacle-avoidance mode, the difference $s-\tau$ equals zero and keeps increasing until the robot switches to the move-to-target mode. This, according to (45) and (48), causes the value of $\lambda$ to go from 00 to $1$ in $t_{s}$ seconds [^2] The parameter $t_{s}$ should be set to a relatively small positive value to ensure that, depending on the current mode of operation, the function $\lambda$ reaches a constant value of either 00 or $1$ before the state $(\xi,\tau)$ enters the jump set of the respective mode.. Similarly, according to (34) and (46), when $\xi\in\mathcal{J}_{1}$, the variables $s$ and $\tau$ are updated to $s+\tau_{s}$ and $s+\tau_{s}+t_{s}$, respectively. Addtionally, during the flow we have $\dot{s}=1$ and $\dot{\tau}=0$. Therefore, as the robot enters in the move-to-target mode, the difference $\tau-s$ equals $t_{s}$ and keeps decreasing until the robot switches to the obstacle-avoidance mode. This, according to (45) and (48), causes the value of $\lambda$ to go from $1$ to 00 in $t_{s}$ seconds. Therefore, by replacing $m$ in (12a) with $\lambda(m,s,\tau)$ one can ensure that the control input vector changes its value between $-\kappa_{s}\mathbf{x}$ and $\kappa_{r}\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$ in a continuous manner, depending on the current mode of operation. The smoothed version of the hybrid control input is then given by
$\mathbf{u}_{sm}(\xi,\tau)=-\kappa_{s}(1-\lambda(m,s,\tau))\mathbf{x}+\kappa_{r}\lambda(m,s,\tau)\mathbf{v}.$ (49)
The resulting hybrid closed-loop system is represented by
$\underbrace{\begin{matrix}\mathbf{\dot{x}}\\ \mathbf{\dot{h}}\\ \mathbf{\dot{a}}\\ \dot{m}\\ \dot{s}\\ \dot{\tau}\end{matrix}\begin{matrix}[l]\;=\mathbf{u}_{sm}(\xi,\tau)\\ \;=\mathbf{0}\\ \;=\mathbf{0}\\ \;=0\\ \;=1\\ \;=0\end{matrix}}_{(\dot{\xi},\dot{\tau})=\mathbf{F}_{sm}(\xi,\tau),{(\xi,\tau)\in\mathcal{F}\times\mathbb{R}}}\quad\quad\underbrace{\begin{matrix}\mathbf{x}^{+}\\ \begin{bmatrix}\mathbf{h}^{+}\\ \mathbf{a}^{+}\\ m^{+}\\ s^{+}\end{bmatrix}\\ \tau^{+}\end{matrix}\begin{matrix}[l]=\mathbf{x}\vspace{0.56cm}\\ \vspace{0.7cm}\in\mathbf{L}(\xi)\\ =\mathbf{U}(\xi)\end{matrix}}_{(\xi^{+},\tau^{+})\in\mathbf{J}_{sm}(\xi),{(\xi,\tau)\in\mathcal{J}_{sm}\times\mathbb{R}}}$ (50)
where $\mathcal{J}_{sm}=\mathcal{J}_{sm}^{0}\cup\mathcal{J}_{sm}^{1}$. Notice that the variable $\tau$ does not change value during the flow, and is updated according to $\mathbf{U}(\xi)$ (46) only when $\xi\in\mathcal{J}_{sm}$.
Remark 5.
The control input vector $\mathbf{u}_{sm}$ allows one to generate smooth robot trajectories that converge to the target location at the origin as long as the robot’s location $\mathbf{x}$ and the mode indicator $m$ are not initialized as follows:
  • $\mathbf{x}(0,0)\in\partial\mathcal{W}_{r_{a}}\cap\mathcal{F}_{1}^{\mathcal{W}}$ and $m(0,0)=0$. In this case, one cannot ensure the forward invariance of the set $\mathcal{K}\times\mathbb{R}$ as $\mathbf{u}_{sm}(\xi(0,0),\tau(0,0))$ at $\mathbf{x}(0,0)$ is directed towards the interior of set $\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}})$.
  • $\mathbf{x}(0,0)\in\mathcal{F}_{0}^{\mathcal{W}}\setminus\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{\mathcal{W}})$ and $m(0,0)=1$. In this case, the obstacle-avoidance control input vector $\mathbf{v}(\mathbf{x},\mathbf{h},\mathbf{a})$ is not viable as the uniqueness of $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ is not guaranteed.
To avoid such situations, one should initialize the robot’s location $\mathbf{x}$ in the interior of the obstacle-free workspace $\mathcal{W}_{r_{a}}$ and ensure that the robot starts operating in the move-to-target mode with the stabilization control input vector $-\kappa_{s}\mathbf{x}$. This can be achieved by setting $m(0,0)=0$ and choosing $\tau(0,0)\in(-\infty,s(0,0))$, as per (45) and (49).
@@ VII-A Control-input smoothing mechanism for practical implementations
2:  Initialize $\mathbf{x}(0,0)\in\mathcal{W}_{r_{a}}$, $\mathbf{h}(0,0)\in\mathcal{W}_{r_{a}}$, $\mathbf{a}(0,0)\in\mathbb{S}^{n-1}$, $m(0,0)\in\mathbb{M}$ and $s(0,0)\in\mathbb{R}_{\geq 0}$. Choose a sufficiently small value of $\bar{\epsilon}$ according to Lemma 2, and initialize $\epsilon\in(0,\bar{\epsilon}].$ Select a minimum safety distance $r_{s}$ such that $r_{s}\in(0,\delta)$ and set the parameters $\gamma,\gamma_{s}$ and $\gamma_{a}$ such that $0<\gamma_{a}<\gamma_{s}<\gamma<\delta-r_{s}$. Choose $R_{s}>r_{a}+\gamma$, used in (8). The parameter $\delta_{s}$, used in (31), is set such that $\delta_{s}\in(0,\tau_{s})$, where $\tau_{s}>0$.
3:  Measure $\mathbf{x}$ and the set ð$\eth_{\mathcal{W}}^{\mathbf{x}}$ as defined in (8).
@@
5:     if $d(\mathbf{x},\eth_{\mathcal{W}}^{\mathbf{x}})\leq r_{a}+\gamma_{s}$then
6:        Identify the set ð$\eth_{i}^{\mathbf{x}}\subset\eth_{\mathcal{W}}^{\mathbf{x}}$ as defined in (17).
7:        if ð$\eth_{i}^{\mathbf{x}}\cap\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})\neq\emptyset$then
@@
16:        if $d(\mathbf{x},\eth_{\mathcal{W}}^{\mathbf{x}})\leq r_{a}+{\gamma},$ then
17:           Identify the set ð$\eth_{i}^{\mathbf{x}}\subset\eth_{\mathcal{W}}^{\mathbf{x}}$ as defined in (17).
18:           if ð$\eth_{i}^{\mathbf{x}}\cap\mathcal{C}_{\mathcal{L}}(\mathbf{x},2r_{a})=\emptyset$then
@@
VIII Simulation Results
We consider an unbounded workspace i.e., obstacle $\mathcal{O}_{0}=\emptyset$, with $2$ three-dimensional, convex obstacles, as shown in Fig. 8. We apply the proposed hybrid feedback controller (12) for the robot initialized at $8$ different locations in the obstacle-free workspace. The target is located at the origin. The radius of the robot is set to $0.2m$. The minimum safety distance $r_{s}=0.1m$ and the parameter $\gamma=0.5m$. The gains $\kappa_{s}$ and $\kappa_{r}$, used in (12a), are set to $1$ and $2$, respectively. The sensing radius $R_{s}$, used in (8), is set to $2m$. The parameter $\epsilon$, used in (28), is set to $0.5m$. From Fig. 8, it can be observed that the robot converges to the target location while simultaneously avoiding collisions with the obstacles. Fig. 9 shows that the center of the robot stays at least $r_{a}$ meters away from the boundary of the obstacles.
Figure 8: Robot trajectories starting from different locations.
Figure 9: Distance of the center of the robot $\mathbf{x}$ from the boundary of the obstacle occupied workspace $\mathcal{O}_{\mathcal{W}}$.
We now consider a two-dimensional workspace containing 12 convex obstacles. A point robot is initialized at 24 different locations along the workspace boundary, as depicted in Fig. 10, with the target located at the origin. The safety distance $r_{s}$ is set to $0.1\,m$. The parameters $\gamma$ and $\epsilon$ are set to $0.2\,m$ and $0.5\,m$, respectively. The gain parameters are $\kappa_{s}=1$ and $\kappa_{r}=2$, and the sensing radius $R_{s}$ is set to $1\,m$. In Fig. 10, dashed trajectories represent the robot’s motion in the move-to-target mode, while solid trajectories indicate motion in the obstacle-avoidance mode. Under the hybrid feedback controller (12), the robot safely converges to the target from all initial locations.
Figure 10: Safe convergence of robot trajectories to the target location.
Next, we examine the effect of varying the sensing radius $R_{s}$ on the robot’s trajectories. The robot with radius $r=0.15m$ is initialized at $[0,9.5]^{\top}$ in a two-dimensional workspace with a single convex obstacle, as shown in Fig. 11. The target location is set at the origin. The safety distance $r_{s}$ is set to $0.05m$. The parameters $\gamma$ and $\epsilon$ are set to $1.5m$ and $0.5m$, respectively. The gains are $\kappa_{s}=1$ and $\kappa_{r}=2$. In Fig. 11, the blue portions of the trajectories correspond to the motion of the robot in move-to-target mode, while red portions correspond to the motion in the obstacle-avoidance mode. When the sensing radius is relatively small, the robot cannot detect all points on the boundary of obstacle $\mathcal{O}_{i}$ that are within its line of sight but outside the sensing region $\mathcal{B}_{R_{s}}(\mathbf{x})$. As a result, the robot has only partial information about the obstacle’s boundary. Consequently, the robot may switch to the move-to-target mode even when it does not have a clear line of sight to the target, as shown in Figs. 11a, 11b, and 11c.
Figure 11: Robot trajectories for different values of the sensing radius $R_{s}$: (a) $R_{s}=1.7m$, (b) $R_{s}=2.5m$, (c) $R_{s}=3.4m$, (d) $R_{s}=4m$.
We next provide a comparison with the separating hyperplane approach developed in [15]. Similar to our approach, this approach can be implemented in a priori unknown environments using the information obtained via a range-bearing sensor mounted on the robot. Contrary to our approach, this approach only works for convex obstacles that satisfy the obstacle curvature condition [15, Assumption 2]. When the workspace consists of obstacles that do not satisfy the obstacle curvature condition, the separating hyperplane approach generates trajectories that converge to an undesired equilibrium (local minimum), as shown in Fig. 12b. On the other hand, the proposed hybrid feedback controller guarantees safe, global asymptotic convergence to the target location, as seen in Fig. 12a. The complete simulation video can be found at333[Online]. Available: https://youtu.be/R6OowaF6dTc.